. 20 2g Fe2O₃ + A, A , ZO3 tre if 20g of fez Oz reacts with ag of AL metal, How many grams or fe could be produced ?
Answer #1

Molar mass of Fe2O3,

MM = 2*MM(Fe) + 3*MM(O)

= 2*55.85 + 3*16.0

= 159.7 g/mol

mass(Fe2O3)= 20.0 g

use:

number of mol of Fe2O3,

n = mass of Fe2O3/molar mass of Fe2O3

=(20 g)/(1.597*10^2 g/mol)

= 0.1252 mol

Molar mass of Al = 26.98 g/mol

mass(Al)= 2.0 g

use:

number of mol of Al,

n = mass of Al/molar mass of Al

=(2 g)/(26.98 g/mol)

= 7.413*10^-2 mol

Balanced chemical equation is:

Fe2O3 + 2 Al ---> 2 Fe + Al2O3

1 mol of Fe2O3 reacts with 2 mol of Al

for 0.1252 mol of Fe2O3, 0.2505 mol of Al is required

But we have 7.413*10^-2 mol of Al

so, Al is limiting reagent

we will use Al in further calculation

Molar mass of Fe = 55.85 g/mol

According to balanced equation

mol of Fe formed = (2/2)* moles of Al

= (2/2)*7.413*10^-2

= 7.413*10^-2 mol

use:

mass of Fe = number of mol * molar mass

= 7.413*10^-2*55.85

= 4.14 g

Answer: 4.14 g